‎‏a ) Describe the transformations to the graph of y = x² to obtain y = -2 ( x + 5 )²- 3 .
b ) Graph y = x² . Then apply the transformations in part ( a ) to graph
y = -2 ( x + 5 ) ² - 3 .
c ) Determine the domain and range of this transformed function .
1
Expert's answer
2020-12-08T06:59:25-0500

y = x² to obtain y = -2 ( x + 5 )²- 3 

= (x − 6)(x − 4)


When x = 0, = 24,

so the y-intercept is 24.


When = 0, x = 4 or x = 6,

so the x-intercepts are 4 and 6.


Taking the average of the x-intercepts,

= 5 is the axis of symmetry.


When x = 5, = (5 − 6)(5 − 4) = −1,

so the vertex is (5, −1).

= − x2 + x + 6


This is an upside-down parabola.

When x = 0, = 6, so this is the y-intercept.

y= − x2 + x + 6= −(x2 − x − 6)= −(− 3)(+ 2)

So the x-intercepts are 3 and −2.


Taking their average, the axis of symmetris x = .


When x = , = − +  + 6 = 6 so the vertex is , 6.

= 5x2 − 20+ 15


When x = 0 then = 15.

y= 5 x2 − 20 x + 15= 5 (x2 − 4+ 3 )= 5 (− 3)( x − 1)

So the two x-intercepts are x = 1 and x = 3.


Hence, the axis of symmetris x = 2 and the vertex is (2, −5).


= (x − 6)(x − 4)


When x = 0, = 24,

so the y-intercept is 24.


When = 0, x = 4 or x = 6,

so the x-intercepts are 4 and 6.


Taking the average of the x-intercepts,

= 5 is the axis of symmetry.


When x = 5, = (5 − 6)(5 − 4) = −1,

so the vertex is (5, −1).

= − x2 + x + 6


This is an upside-down parabola.

When x = 0, = 6, so this is the y-intercept.

y= − x2 + x + 6= −(x2 − x − 6)= −(− 3)(+ 2)

So the x-intercepts are 3 and −2.


Taking their average, the axis of symmetris x = .


When x = , = − +  + 6 = 6 so the vertex is , 6.

= 5x2 − 20+ 15


When x = 0 then = 15.

y= 5 x2 − 20 x + 15= 5 (x2 − 4+ 3 )= 5 (− 3)( x − 1)

So the two x-intercepts are x = 1 and x = 3.


Hence, the axis of symmetris x = 2 and the vertex is (2, −5).